An item sells with a markup of 37% on selling price. What is the markup as a percentage on cost?
Answer #1

Let sales price = 100

Markup = 100*37% = 37

Cost = 100-37 = 63

Markup as a percentage on cost = Markup/salea

= 37/63

= 58.73%